Circular No. 7103

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J1723–376 F. E. Marshall, Goddard Space Flight Center (GSFC); and C. B.

Markwardt, National Academy of Sciences/National Research Council and GSFC, on behalf of the RXTE Science Operations Center, report the discovery of a transient x-ray source with quasiperiodic oscillations (QPO) at 816 Hz with an rms amplitude of 5 percent. The transient was first seen on Jan. 25.21 UT during a scan with the PCA on RXTE, with a 2–10-keV flux of 1.5 × 10^{−9} erg cm^{−2} s^{−1}. The QPO were detected during a pointed observation at Jan. 31.1, when the flux was 1.2 × 10^{−9} erg cm^{−2} s^{−1}. The best-fit source position is α = 17^{h}23^{m}.6, δ = −37^{o}39′ (equinox 2000.0), with an estimated uncertainty of 2′. Additional RXTE observations are planned.

AB AURIGAE N. M. Ashok and T. Chandrasekhar, Physical Research Laboratory

(PRL), Ahmedabad; and H. C. Bhatt and P. Manoj, Indian Institute of Astrophysics (IIA), Bangalore, report: “AB Aur (HD 31293), the prototype and one of the brightest Herbig Ae/Be stars, presently seems to have entered an unusually inactive phase. A near-infrared spectrum taken with the 1.2-m telescope (+ Nearinfrared Imager/Spectrometer + 256×256 HgCdTe NICMOS3 array) at the PRL’s Mt. Abu Observatory on Jan. 9 shows that the Paβ line is marginally in absorption and the Brγ emission is absent. Optical linear polarization measurements of AB Aur, made with a polarimeter on the IIA 1-m telescope at Kavalur on Jan. 15, show that its degree of polarization (P (V) = 0.8 ± 0.1 percent, θ = 70^{o} ± 3^{o}) is now at its highest. The photometric magnitudes were V = 7.0 ± 0.1 on Jan. 15; and J = 7.2 ± 0.15, H = 6.4 ± 0.15, K = 6.0 ± 0.15 on Jan. 9. While its visual magnitude is at its brightest, the infrared J, H, K magnitudes are at their faintest level, indicating substantial decrease in the near-infrared caused by circumstellar dust.”

SUPERNOVAE 1998S, 1998aq, 1998bu, 1998es, 1999D, 1999E Unfiltered CCD magnitude estimates by L. Kiss, K. Sárneczky, and

E. Barát, Piszkéstető, Konkoly Observatory: SN 1998S in NGC 3877, Jan. 23.138 UT, 18.0 ± 0.4; SN 1998aq in NGC 3982, Jan. 21.999, 19.6 ± 0.4; SN 1998bu in NGC 3368, Jan. 21.985, 17.7 ± 0.3; SN 1998es in NGC 632, Jan. 21.706, 15.8 ± 0.2; SN 1999D in NGC 3690, Jan. 21.992, 16.6 ± 0.3; SN 1999E, Jan. 23.156, 16.7 ± 0.3.

1999 February 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green